OpenStudy (anonymous):

Help! Medal + fan! How many solutions will this system of equations have? y = -3.5x - 3.5 y = -3.5x + 3.5

OpenStudy (anonymous):

0. The lines have the same slope, but a different constant value. So they will logically never touch.

OpenStudy (anonymous):

o thank u so much do u think u could help me on another one please

OpenStudy (anonymous):

Given the equation 5x + y = 7, which equation below would cause a consistent-independent system? 10x + 2y = 14 -15x - 3y = -6 5x + y = -7 6x + 2y = 7

OpenStudy (anonymous):

@IAmSinged

OpenStudy (anonymous):

Well, I'd start by reordering the equations into y = mx + b form. 2y = -10x + 14 -3y = 15x + 6 y = -5x + 7 2y = -6x + 7 Then we divide the value in front of the x term by the value in front of the y term (the slope!). -5 -5 -5 -3. Doing the same for the original equation we get -5. If the slopes are the same that means 1 of 2 things: 1. They're the same equations 2. They're parallel and will never touch So the last option is right.

OpenStudy (anonymous):

thank u sooo much u really helped a lot